OpenStudy (albert0898):

Find four consecutive odd integers such that five times the sum of the first and third integers exceeds four times the sum of the second and last integers by 14.

OpenStudy (albert0898):

The answer is not -5, -3, or -1. The answer is 13, 15, 17, 19.

OpenStudy (albert0898):

I updated the question!!!!

OpenStudy (radar):

2x+1 = 1st odd integer 2x+3 = 2nd 2x+5 =3rd 2x+7 = 4th 5(2x+1+2x+5)= 20x+30) is the sum of the 1st & 3rd times 5 4(2x+3+2x+7) =16x+40) is the sum of 2nd & last times 4 20x+30 = 16x + 40 +14 4x = 24 x=6 1st integer is 13 2nd is 15 3rd us 17 4th is 19 Check out: 5 (13 +17) = 4(15+19)+14 150 =150

OpenStudy (radar):

You knew the answer, did you follow the procedure for solution with understanding?

OpenStudy (albert0898):

Why'd you do 2x+1 = 1st odd integer 2x+3 = 2nd 2x+5 =3rd 2x+7 = 4th Why not just do: x, x + 3, x + 5, x +7

OpenStudy (radar):

I wanted to insure my results were odd, in case there were even digits that would result in sums differing by 14.

OpenStudy (albert0898):

So you could you still do x, x + 2, and so on, or no?

OpenStudy (radar):

Have you tried it using x, and x+2, etc. It may work just as well.

OpenStudy (radar):

I can try if you haven't.

OpenStudy (albert0898):

It comes to the same thing. But may you please explain to me why the term "exceeds four times... by 14." I thought this meant to add 14 to 10x + 20, but it means to add 14 to 8x + 46.

OpenStudy (radar):

x = 1st digit x+2 =2nd x+4 =3rd x+6 = 4th 5(x+x+4)=4(x+2+x+6)+14 10x+20=8x+32+14 2x=26 x=13 Yes it works so there are apparently only those numbers that will work, it is easier to do it that way, but I always make sure I have a odd result when looking for odd. lol.

OpenStudy (radar):

It means that the first sum is larger by 14 than the second sum. So to make them equal you add 14 to the second sum.

OpenStudy (albert0898):

Oh, I see now. Thank you for your help! xD

OpenStudy (radar):

You know when they are asking about even numbers I use 2x to be sure I get even numberts, I guess I am sort of paranoid about numbers.

OpenStudy (radar):

Good luck with these.

OpenStudy (radar):

You are most welcome.